A water tank has 8 feet of water in it. At the end of each hour the water level falls to 75% of the previous level. Identify the graph of the sequence and the correct description of its pattern. Then identify the what the water level will be after 6 hours.
1
Expert's answer
2019-07-04T10:20:50-0400

1) Let An is the level of water after n haours. Then:

A0 = 8 feet,

A1 = A0 * 75/100 = 8 * 3/4 = 6;

A2 = A1 * 3/4 = 6 * 3/4 = 9/2 = 4 1/2;

A3 = A2 * 3/4 = 9/2 * 3/4 = 27/8 = 3 3/8;

A4 = A3 * 3/4 = 27/8 * 3/4 = 81/32 = 2 17/32;

A5 = A4 * 3/4 = 81/32 * 3/4 = 243/128 = 1 115/128;

A6 = A5 * 3/4 = 243/128 * 3/4 = 729/512 = 1 217/512 (feet).


2) The common formula for the water level after n haours is:

An = 8 * (3/4)n.

This is the exponential decay with the base 3/4. All the points on a line that graphs the pattern are accurate:



3) Thus, the water level after 6 hours will be:

 A6 ≈ 1,423 feet.